Cati's Catering is planning an office reception. The office administrator has requested a candy mixture that is 25% chocolate. Cati has available mixtures that are either 50% chocolate or 10% chocolate. How much of each type should be mixed to get a 20-lb mixture that is 25% chocolate?
Answer by stanbon(75887)

Cati's Catering is planning an office reception. The office administrator has requested a candy mixture that is 25% chocolate. Cati has available mixtures that are either 50% chocolate or 10% chocolate. How much of each type should be mixed to get a 20-lb mixture that is 25% chocolate?
-------------------------
Equation:
chocolate + chocolate = chocolate
0.50x + 0.10(20-x) = 0.25(20)
Multiply thru by 100 to get:
50x + 10*20 - 10x = 25*20
40x = 5*20
x = 2.5 lb (amt of 50% chocolate needed in the mixture)
---
20-2.5 = 17.5 lb (amt of 10% chocolate needed in the mixture)
=======================
Cheers,
Stan H.